OpenStudy (anonymous):

Gabriel measured the length of two icicles hanging from his porch. The first icicle was 27 centimeters 3 millimeters and the second icicle was 11 centimeters 7 millimeters. How much longer is the first icicle than the second icicle?

OpenStudy (anonymous):

these are the answers. 15 centimeters 6 millimeters 16 centimeters 4 millimeters 16 centimeters 6 millimeters

OpenStudy (johnweldon1993):

to solve this I'll justt convert the cm to mm.. there are 10 mm in 1 cm if I'm not mistaken so first icicle 27 cm = 27(10)mm = 270 mm....+ 3 mm more is 273 mm total... 11cm = 11(10)mm = 110mm..... + 7mm more is 117 mm total so now subtract them 273mm - 117mm is 156mm now to convert that to cm....divide by 10 156 mm = 156/10 cm = 15 cm and 6 mm

OpenStudy (anonymous):

thanks help me with others plzzzz